2023 July

From EOVSA Wiki
Jump to navigation Jump to search

Back to Observing Log

July 16

23:00 UT In the afternoon all of the Laird temperature controllers stopped regulating, allowing the FEMs to reach >40 C. This was likely heat-related due to unusually high ambient temperatures. Things were back to normal by the next day.

July 22

00:10 UT The new feed + radome was installed on Ant 11 and the control files were changed to add Ant 11 tracking. The pointing measurement at 21:30 UT shows that Ant 11 is not pointing well in HA, so this needs to be updated. I will take more measurements tomorrow and adjust the pointing. The phase calibration at 23:14 UT showed coherence at low frequencies on Ant 11, so I was able to adjust the delay (about 3 ns) to optimize that. We have a weekend coming up, so we won't be able to explore focus adjustments, but I will do a series of observational tests with this current focus over the weekend.

July 24

17:23 UT The feed orientation was off by 45 degrees. Owen traced that to having the plate installed backwards, so it is now flipped and the orientation should be right. I am doing a SOLPNTCAL to check it.

23:06 UT Still no good. I went through the calculation and pointed out what is needed, but it means machining new slots in the plate, so there will be a delay of a day or so.

July 25

17:00 UT LNA Drain voltages on the 27m were all zero. I restarted the LNA and also rebooted the brick and antenna controller.

July 26

15:00 UT ACC needed to be restarted.

July 28

17:22 UT ACC locked up again. I perfomed a full power cycle.